# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v2

@@ 2.1 Conditions for Point Identification
Blue lines represent the distribution of $Y(1)$ conditional on a given value of $X=x$, and red lines the distribution of $Y(0)$. Panels in the first row show the cumulative distributions, and in the second row, the probability density functions. The left column shows a scenario with partial identification, and the right column shows a scenario with point identification. The green dashed line represents the point $t$ that maximizes $F_{1}(t|x)-F_{0}(t|x)$. In both scenarios $\min_{t}F_{1}(t|x)-F_{0}(t|x)=0$, for example for $t=0$.
@@ 4 Large Sample Inference with Cross-Fitting
The approach of the previous section uses sample splitting to provide a strong guarantee of coverage under weak conditions. In this section, I propose a new approach that uses cross-fitting, a form of sample splitting that uses the whole sample for estimating both the covariate-adjustment term and the final estimator, in order to improve power (for a discussion on cross-fitting see, for example, Chernozhukov et al., 2018).
In order to derive asymptotically valid confidence intervals, I define new regularity conditions (A2) and show in Thorem 3 an intermediate result, that the estimators are asymptotically Gaussian. The main result is then established in Theorem 4, which shows uniform validity of CIs for $\theta$. Finally, Theorem 5 shows conditions under which the CIs are exact.
@@
A2(i) restricts the possibility of point masses in the distribution of the outcome. A2(ii) simply requires the estimators $(\hat{s}_{L},\hat{s}_{U})$ to have any limit in probability, at any rate of convergence, and allows them to be inconsistent for the sharp $(s^{*}_{L,P},s^{*}_{U,P})$. A2(iii) is assumed to simplify exposition. It ensures the asymptotic distribution is first-order insensitive to the choice of the optimizers in Definition 2(iii), and it is a generalization for $P\in\mathcal{P}$ of the assumption used in Fan and Park (2010) in their approach to inference on Makarov bounds without covariates. I discuss two approaches for relaxing A2(iii) in the presence of multiple optimizers in the Online Appendix (OA.3). One of them consists in using the bootstrap for inference instead of relying on normal approximations to $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$. Example 1 illustrates a data generating process where A2(iii) holds. A2(iv) is used to ensure that the problem is not trivial in the sense that $Var[\sqrt{n}\widehat{\theta}_{L}]$ and $Var[\sqrt{n}\widehat{\theta}_{U}]$ do not converge to zero. It is only used to apply the result from Stoye (2009) to get asymptotic validity of the two-sided CI in Theorem 4.
@@ 7 Conclusion
I illustrate the practical relevance of the method in a simulation study and an application to microcredit. The application reveals an important presence of heterogeneity, with significant negative and positive treatment effects. The results show that the method is able to provide informative bounds even in challenging contexts, such as when the ATE is not statistically significant, the sample size is small, and the number of covariates is large.
By using CAIDE, an applied researcher can learn more about heterogeneous effects not only by gathering larger samples but also by collecting information on covariates that help predict potential outcomes.
@@ A.2 Estimated vs True Propensity Score
The result that using the estimated propensity score yields smaller asymptotic variance than using the true propensity score is similar to other results in the literature (e.g., Hirano et al., 2003 in the context of the ATE).
A.3 Estimation of the covariate-adjustment term
@@
A.4 Definitions Ommited from the Main Text
@@
A.5 Proofs of the Main Results
@@ A Taylor expansion gives
Supplemental Appendix to “Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach”
@@
    Note that, as before, the elements of the variance-covariance matrix can be estimated using sample analogues.

### v3

@@ 2.1 Conditions for Point Identification
Blue lines represent the distribution of $Y(1)$ conditional on a given value of $X=x$, and red lines the conditional distribution of $Y(0)$. Panels in the first row show the cumulative distributions, and in the second row, the probability density functions. The left column shows a scenario with partial identification, and the right column shows a scenario with point identification. The green dashed line represents the point $t$ that maximizes $F_{1}(t|x)-F_{0}(t|x)$. In both scenarios $\min_{t}F_{1}(t|x)-F_{0}(t|x)=0$, for example for $t=0$.
@@ 4 Large Sample Inference with Cross-Fitting
The approach of the previous section uses sample-splitting to provide a strong guarantee of coverage under weak conditions. In this section, I propose a new approach that uses cross-fitting to improve power. Unlike sample-splitting, cross-fitting uses the whole sample for estimating both the covariate-adjustment term and the final estimator.
In order to derive asymptotically valid confidence intervals, I define new regularity conditions (A2) and show in Theorem 3 an intermediate result, that the estimators are asymptotically Gaussian. The main result is then established in Theorem 4, which shows uniform validity of CIs for $\theta$. Finally, Theorem 5 shows conditions under which the CIs are exact.
@@
A2(i) restricts the possibility of point masses in the distribution of the outcome. A2(ii) requires the estimators $(\hat{s}_{L},\hat{s}_{U})$ to have any limit in probability, at any rate of convergence, and allows them to be inconsistent for the sharp $(s^{*}_{L,P},s^{*}_{U,P})$. A2(iii) is assumed to simplify exposition. It ensures the asymptotic distribution is first-order insensitive to the choice of the optimizers in Definition 2(iii), and it is a generalization for $P\in\mathcal{P}$ of the assumption used in Fan and Park (2010) in their approach to inference on Makarov bounds without covariates. I discuss two approaches for relaxing A2(iii) in the presence of multiple optimizers in the Online Appendix (OA.3). One of them allows dropping A2(iii) by using the bootstrap for inference instead of relying on normal approximations to $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$. Example 1 illustrates a data generating process where A2(iii) holds. A2(iv) is used to ensure that the problem is not trivial in the sense that $Var[\sqrt{n}\widehat{\theta}_{L}]$ and $Var[\sqrt{n}\widehat{\theta}_{U}]$ do not converge to zero. It is only used to apply the result from Stoye (2009) to get asymptotic validity of the two-sided CI in Theorem 4.
@@ 7 Conclusion
I illustrate the practical relevance of the method in a simulation study and an application to microcredit. The application reveals an important presence of heterogeneity, with statistically significant negative and positive treatment effects. The results show that the method is able to provide informative bounds even in challenging contexts, such as when the ATE is not statistically significant, the sample size is small, and the number of covariates is large.
By using CAIDE, an applied researcher can learn more about heterogeneous effects not only by gathering larger samples but also by collecting information on variables that help predict potential outcomes.
@@ A.2 Estimated vs True Propensity Score
A.3 Asymptotic Distribution of Plug-in Estimator Based on Fan and Park (2010)
I argue that it is generally difficult to characterize the asymptotic distribution of a plug-in estimator based on Fan and Park (2010)’s representation of the identified set (2.5) when the conditional distributions are estimated nonparametrically. I show that the asymptotic distribution depends on the rate of convergence of the estimators of $F_{1}(t|x)$ and $F_{0}(t|x)$, and that different models/algorithms used to estimate these functions lead to different asymptotic distributions of the plug-in estimator. This is true even when using sample-splitting. For simplicity, I focus on the lower bound.
Define
$\Delta(t|x)=F_{1}(t|x)-F_{0}(t|x).$
Let $\hat{F}_{1}(t|x)$ and $\hat{F}_{0}(t|x)$ be estimators respectively of $F_{1}(t|x)$ and $F_{0}(t|x)$, and $\hat{\Delta}(t|x)=\hat{F}_{1}(t|x)-\hat{F}_{0}(t|x)$. Define $\hat{s}(x)\in\arg\min_{t}\hat{\Delta}(t|x)$. Since my estimation strategies in Sections 3 and 4 use sample-splitting, I also use sample-splitting to define the plug-in estimator based on (2.5). Consider a random split of $\{1,\dots,n\}$ into three equal-sized sets $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ ($n$ is assumed a multiple of 3 for simplicity). Let $\hat{\Delta}$ be estimated using data from $\mathcal{I}_{1}$ only, and $\hat{s}$ estimated using data from $\mathcal{I}_{2}$ only. The plug-in estimator of $\theta_{L}^{*}$ based on (2.5) is given by
$\hat{\theta}_{FP}=\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i}).$
Let $\mathcal{D}_{j}=(Y_{i},D_{i},X_{i})_{i\in\mathcal{I}_{j}}$ for $j=1,2,3$. Define
$\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})=\mathbb{E}[\hat{\Delta}(\hat{s}(X)|X)|\mathcal{D}_{1},\mathcal{D}_{2}],$
where $X$ is independent of $\mathcal{D}_{1},\mathcal{D}_{2}$. It follows that
$\left(\hat{\theta}_{FP}-\theta_{L}^{*}\right)=\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)+\left(\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}\right).$
Since the first term is mean-zero, it follows that
$\sqrt{n}\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)$
is asymptotically normal by a triangular array CLT. The second term, $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$, is random since the data $(\mathcal{D}_{1},\mathcal{D}_{2})$ is random. $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ may or may not be bounded in probability when multiplied by $\sqrt{n}$. In general, the rate of convergence of this term depends on the rate of convergence of $\hat{\Delta}$. Since $\hat{\Delta}$ is estimated nonparametrically, this rate may be slow when regularity conditions such as sparsity and specific forms of smoothness don’t hold. In contrast, the asymptotic distribution in Theorem 3 does not rely on such conditions. The asymptotic distribution of an appropriately scaled version of $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ depends on which model/algorithm is used to estimate $\hat{\Delta}$, since different estimators $\hat{\Delta}$ lead to different values of $\bar{\Delta}$. For example, the distribution of $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})$ is different and difficult to characterize if $\hat{\Delta}$ is calculated with random forests or neural networks, which are some of the models I use in Section 6.
Note that the term $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ may be positive or negative. If $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})$ is symmetrically distributed around $\theta_{L}^{*}$, $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}>0$ with probability $1/2$. In this case, a hypothesis test for $H_{0}:\theta_{L}^{*}\leq\tau$, $\tau\in[0,1]$, based on the asymptotic distribution of $\sqrt{n}\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{F}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)$ may overreject. That is, if one conducts a hypothesis test with significance level $\alpha$ based on the incorrect assumption that $\sqrt{n}\left(\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}\right)=o_{P}(1)$, the test may, for example, reject $\theta_{L}^{*}=0$ when $\theta=0$ with probability higher than $\alpha$.
In contrast, my approaches to inference do not depend on the rate of convergence or the specific model/algorithm used to estimate $s^{*}_{L}$. Theorem 2 provides a finite-sample guarantee of coverage under no restrictions on how $\hat{s}_{L}$ is calculated. In Theorem 4, I provide a large-sample guarantee of coverage under no restrictions on the rate of convergence of $\hat{s}_{L}$ or which model/algorithm is used to calculate it. The result of Theorem 4 is possible because of the alternative characterization of the identified set in Theorem 1. Unlike (2.5), which relies on the average of conditional expectations, (2.3) is expressed in terms of unconditional expectations. As a consequence, Theorem 3 shows that estimation error in $\hat{s}_{L}$ translates into negative bias,
$\mathbb{E}[\sqrt{n}(\widehat{\theta}_{L}-\theta_{L}^{*})]\leq 0+o_{P}(1),$
which makes the estimate $\widehat{\theta}_{L}$ conservative. Since $\theta_{L}^{*}$ is a lower bound, an estimate $\widehat{\theta}_{L}$ that is too low leads to a confidence set that may be conservative, but valid.
A.4 Estimation of the covariate-adjustment term
@@
A.5 Definitions Ommited from the Main Text
@@
A.6 Proofs of the Main Results
@@ A Taylor expansion gives
Supplemental Appendix to “Predicting the Distribution of Treatment Effects via Covariate-Adjustment,
with an Application to Microcredit”
@@
    As before, the covariance matrix can be estimated with sample analogues.